OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER!!! What is the solution to g over five equals three over two? 7.5 4 3 10

Mehek (mehek14):

\(\dfrac{g}{5}=\dfrac{3}{2}\) like that?

OpenStudy (iiariana):

Convert that into letters and numbers. g5=32

OpenStudy (iiariana):

Do this: Multiply both sides by 5. Multiply both sides by 2. See what comes up.

Mehek (mehek14):

to solve for g, multiply \(\dfrac{3}{2}*5\)

OpenStudy (iiariana):

And if you do that you'll get ?.5

OpenStudy (anonymous):

7.5 Thanks guys!

OpenStudy (iiariana):

Yes correct! And np

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (anonymous):

Wish I could medal u both, but i can. So I will fan both :)

OpenStudy (iiariana):

It's fine :-)

OpenStudy (anonymous):

case closed :) Thanks guys!!!